OpenStudy (anonymous):

help!!!! Write the value for x that makes the following equation true. 7^x+4=7^10

jimthompson5910 (jim_thompson5910):

assuming this is 7^(x+4) = 7^10 since the bases are both equal to 7, this means the exponents must be equal to each other

jimthompson5910 (jim_thompson5910):

so, x+4 = 10 x = ??

OpenStudy (anonymous):

6

jimthompson5910 (jim_thompson5910):

yep